# Graph of AI Ideas: Leveraging Knowledge Graphs and LLMs for AI Research Idea Generation

### v1

@@ Abstract
Reading relevant scientific papers and analyzing research development trends is a critical step in generating new scientific ideas. However, the rapid increase in the volume of research literature and the complex citation relationships make it difficult for researchers to quickly analyze and derive meaningful research trends. The development of large language models (LLMs) has provided a novel approach for automatically summarizing papers and generating innovative research ideas. However, existing paper-based idea generation methods either simply input papers into LLMs via prompts or form logical chains of creative development based on citation relationships, without fully exploiting the semantic information embedded in these citations. Inspired by knowledge graphs and human cognitive processes, we propose a framework called the Graph of AI Ideas (GoAI) for the AI research field, which is dominated by open-access papers. This framework organizes relevant literature into entities within a knowledge graph and summarizes the semantic information contained in citations into relations within the graph. This organization effectively reflects the relationships between two academic papers and the advancement of the AI research field. Such organization aids LLMs in capturing the current progress of research, thereby enhancing their creativity. Experimental results demonstrate the effectiveness of our approach in generating novel, clear, and effective research ideas.
1 Introduction
The generation of creative ideas is a vital aspect of scientific research, fueling technological innovation and breakthroughs. Traditionally, this process has been human-driven, requiring researchers to consult extensive literature, summarize research trends, and propose improvements or optimizations as new research directions. However, in the rapidly evolving field of artificial intelligence, the increasing volume of literature and the complexity of interdisciplinary research present challenges. Researchers are constrained by the vastness of scientific literature and intricate citation relationships, demanding significant time and effort. Consequently, the overwhelming number of related papers makes this task particularly difficult, highlighting the urgent need for efficient idea generation and innovation.
Recent advancements in large language models (LLMs), such as ChatGPT [^1] https://chatgpt.com/, Claude [^2] https://claude.ai/, and LLaMA Touvron et al. (2023a, b), have demonstrated substantial potential in fostering innovation. Existing studies have shown that leveraging the specialized knowledge and exceptional natural language generation abilities acquired through large-scale pretraining, LLMs are capable not only of understanding and generating complex academic content Si et al. (2024); Kumar et al. (2024) but also excelling in coordinating constructing implicit chains of thought and revealing subtle connections between papers Wang et al. (2024b); Li et al. (2024). Utilizing LLMs to assist researchers in generating new ideas is of significant importance for enhancing research productivity.
However, current methods of idea generation based on LLMs still face three main challenges: 1) When LLMs are tasked with generating ideas, they are provided with a substantial amount of research literature and asked to summarize development trends based on the sequence and citation relationships between papers. However, the semantic information embedded within these citation relationships is not adequately considered. For example, a cited paper could be used as a baseline or as a counterexample, which makes it easy for LLMs to be influenced by less relevant works, potentially leading to ideas that lack logical coherence and technical innovation. 2) The approach of prompting LLMs to analyze citation relationships and establish chains of thought only summarizes development trends in a simple linear structure. This one-dimensional data structure oversimplifies the summary of research trends. In fact, the relationships between academic papers can be represented as a complex graph structure through knowledge graphs. Summarizing and mining more intricate relationships can yield more diverse research ideas. 3) Existing methods for generating research ideas using LLMs typically rely on human evaluation and prompt-based assessments, lacking an objective evaluation mechanism for assessing, providing feedback, and refining the generated ideas.
To address the aforementioned challenges and limitations, this paper proposes a method for generating novel AI research ideas based on knowledge graphs and large language models, specifically aimed at the rapidly evolving field of artificial intelligence, which is characterized by a large volume of open-access papers. The proposed method, Graph of AI Ideas (GoAI), works as follows. First, GoAI searches for relevant literature based on the topic under investigation and selects the most pertinent papers as key references. Once the key references are identified, GoAI analyzes their cited references and the articles that cite them, constructing a knowledge graph, the GoAI Graph, that represents the research trends of the topic. In this knowledge graph, entities represent papers, and the relations between entities reflect the semantic descriptions of citation relationships within the papers, such as “based on…”, “support by…”, “in contrast …”, etc. Furthermore, these relationships also incorporate the position where the references are cited in the original text to assess the importance of different citing papers. After constructing the graph, we employ an agent based on large language models, the GoAI Agent, to collaborate with the knowledge graph. Using a beam search algorithm, the large language model dynamically explores the graph, investigating multiple meaningful development paths that may lead to new research ideas. These paths are explored based on the entities, the semantic relationships, and the importance of these relationships within the graph. The model then generates new research ideas, performs novelty checks, and iteratively refines them. Further evaluation confirms the effectiveness of our method.
The main contributions of this paper are as follows:
  1. 1.
    We propose a novel academic knowledge graph, GoAI, and its construction method, using the AI field as an example. This graph not only represents citation relationships between papers but also captures the semantic information contained within these citations and differentiates the importance of various references.
  2. 2.
    We introduce the GoAI Agent, a research idea generation agent that collaborates with the knowledge graph. By enabling the large language model to dynamically explore the AI idea graph, the agent analyzes trends and generates new ideas based on this exploration.
  3. 3.
    We present an evaluation method for assessing the research ideas generated by large language models for assessing, providing feedback, and refining. Experimental results demonstrate the effectiveness of the GoAI method we propose.
2 Related Works
2.1 AI-based Scientific Discovery
@@
Artificial intelligence has become a powerful tool for advancing research processes, reshaping scientific discovery by reducing the time needed for experimental iterations, and assisting researchers in various capacities Yu et al. (2024); Abramson et al. (2024); de Haan et al. (2024); Tan et al. (2024). The development of large language models, with their advanced natural language processing capabilities and ability to integrate interdisciplinary information, enables more efficient literature analysis and the generation of novel hypotheses Kumar et al. (2024); Si et al. (2024). Prior research has enhanced novelty by iteratively comparing AI-generated ideas with existing literature Wang et al. (2024a), while other studies have focused on linking academic knowledge through core papers, expanding understanding via academic graphs and entity-centered knowledge stores Baek et al. (2024); Wang et al. (2024b); Li et al. (2024). Some studies use multi-agent systems based on LLMs to foster the generation of research ideas in autonomous scientific discovery Su et al. (2024). In this paper, we adopt a paradigm for generating research ideas based on academic graphs derived from core papers. Unlike previous works, we represent citation relationships between academic papers and summarize development trends through a knowledge graph, rather than through research trend chains. Furthermore, we introduce the concept of constructing a knowledge graph that incorporates semantic information and citation importance, providing richer information for analyzing relationships between papers.
2.2 Graph-augmented LLMs
@@
The integration of large language models with knowledge graphs, which store information in natural language, is an emerging research area aimed at fully utilizing the explicit and updatable external knowledge within structured graphs to enhance the capabilities of large language models Pan et al. (2024); Huang et al. (2024), particularly in tasks such as reasoning Luo et al. (2024b, c) and rule completion Luo et al. (2024a) and benchmarking Ruan et al. (2025); Zhu et al. (2024); Ruan et al. (2024b, a) . To input the structured knowledge from knowledge graphs into large language models, some methods transcribe entity-relation-entity triples from the knowledge graph into natural language text that is understandable by the language model, which is then provided as a prompt to the LLM Sun et al. (2024), similar to retrieval-augmented generation. Other methods directly input graph information processed by approaches such as Graph Neural Networks (GNNs) into the large language model, utilizing the implicit graph information for enhancement Huang et al. (2024); Mavromatis and Karypis (2024). In this paper, since the literature used to construct the graph is already in natural language and the relations within the graph contain semantic information in natural language, we adopt the method of transcribing the graph into textual prompts for input into the large language model to achieve the collaboration between the graph and the model.
3 Methodology
@@
Figure 1: The framework of GoAI. The framework consists of four stages: (1) Literature Search and Filtering, (2) GoAI Graph Construction, (3) Idea Generation, and (4) Novelty Evaluation.
3.1 GoAI Framework
In this section, we will provide a detailed description of the GoAI method framework. As shown in Figure 1, the framework consists of four stages: (1) Literature Search and Filtering, (2) GoAI Graph Construction, (3) Idea Generation, and (4) Novelty Evaluation.
First, given a research topic, GoAI searches the online literature database for relevant papers and identifies key reference papers based on their ranking. Based on the citation relationships of these key reference papers, the graph is extended both forward and backward to reflect various research trends in the field. Then, using LLM-as-agent, the knowledge graph is analyzed to predict future research directions. The generated ideas undergo novelty checks via structured-thinking evaluation, with iterative refinement and integration leading to the final ideas. The prompts used in this section can be found in our supplementary materials.
3.2 Literature Search and Filtering
@@
We use the Semantic Scholar API [^3] https://www.semanticscholar.org/product/api to search for papers in the online paper database. Given an initial research topic, we retrieve the top-ranked papers, which we refer to as key references, forming the foundation for building the graph. Starting from the corresponding key reference papers, GoAI constructs the graph by extending in both forward and backward directions based on citation relationships. The backward extension is derived from the reference list of the key reference, tracing the origins of the paper’s ideas. The forward extension is achieved by exploring which papers in the literature cite the key reference paper, representing the subsequent development of the idea.
@@
3.3 GoAI Graph Construction
3.3.1 GoAI Graph Overview
The GoAI graph is used to represent the relationships between papers that have citation relationships, including the citing and cited papers, the importance of the citation (measured by the position where the reference appears in the original text), and the semantic information contained in the citation. The general representation of the GoAI graph is a structured <paper, citation position, citation semantics, paper> quadruple. The citation position and citation semantics together form the relation field in traditional knowledge graphs, which is used for subsequent analysis of the graph. To construct the GoAI graph, it is necessary to separately obtain the papers as entities and the citation positions and citation semantics as relations.
3.3.2 Relation Definition and Extraction
@@
3.4 Idea Generation
The process of generating ideas from the graph is divided into two steps: graph exploration and idea generation. Figure 2 illustrates an example workflow of graph exploration and idea generation.
3.4.1 Graph Exploration
In the graph, the entities are papers, and the relations are represented as pairs of citation positions and citation semantics. At the beginning of the $D$-th iteration, each path consists of $12D-$ 1 quadruples, i.e., $p=\{(e^{d}_{s},r^{d},e^{d}_{o})\}^{2D-1}_{d=1}$, where $e_{s}$ and $e_{o}$ represent the subject and object entities, respectively, and $r$ denotes the specific relation between them. The pairs $(e^{d}_{s},r^{d},e^{d}_{o})$ and $(e^{d+1}_{s},r^{d+1},e^{d+1}_{o})$ are interconnected. The sets of endpoint entities and relations of all paths in the path set $P$ are denoted as $E^{D-1}=\{e^{d-1}_{1},e^{d-1}_{2},\cdots,e^{d-1}_{N}\}$ and $R^{D-1}=\{r^{d-1}_{1},r^{d-1}_{2},\cdots,r^{d-1}_{N}\}$ , respectively.
The exploration phase of the $D$-th iteration aims to utilize the LLM to identify the top-N most relevant entities $E^{D}$ from the neighboring entities of the current top-N endpoint entity set $E^{D-1}$, and use $E^{D}$ to extend the top-N paths in the path set $P$. The endpoint entities here can be either the head or tail entities of the path, corresponding to backward and forward extensions, respectively. To address the complexity of processing a large number of neighboring entities using the LLM, inspired by the approach of ToG Sun et al. (2024), this paper implements a two-step exploration strategy: first, explore the important relations, and then use the selected relations to guide entity exploration.
Relation Exploration The relation exploration is a beam search process with a depth of 1 and a width of $N$, from $E^{D-1}$ to $R^{D}$ The entire process can be broken down into two steps: search and pruning. We use the GoAI Agent, an LLM-based agent, to automatically complete this process.
@@ 3.4.1 Graph Exploration
[…]
@@
3.4.2 Idea Generation with paths
After obtaining the current path $p$ through the exploration process, we prompt the LLM to evaluate the development trends of the papers in the current path, which guides the generation of new ideas. For each generated path, the model predicts the potential future trends. Once the future trends are identified, we continue to prompt the LLM to elaborate on the motivations, novelty, and methodology, ultimately solidifying the idea. Through this gradual process, more detailed ideas can be generated. We also use the novelty checking agent introduced in Section 3.5 to assess candidate ideas and iteratively refine them.
3.5 Novelty Evaluation with GoAI-CoT-Reviewer
Recent research indicates that a significant factor influencing the capabilities of LLMs is the systematicity and structured nature of the reasoning process. Systematicity refers to the model’s ability to perform multi-stage reasoning rather than arriving at a conclusion in a single step. Structuring refers to the model’s ability to clearly identify its current stage of thinking and understand the primary task to be addressed in each stage. OpenAI’s o1 Zhong et al. (2024) effectively addressed these issues by enabling the model to independently perform systematic and structured reasoning through language, while LLaVA-CoT Xu et al. (2025) enhanced the model’s step-by-step reasoning ability via supervised fine-tuning.
Given that the evaluation of paper idea novelty requires the model to deeply analyze the content of the idea and provide a rating, the model needs to engage in profound, step-by-step thinking. Therefore, we evaluate the novelty of generated ideas through structured, step-by-step assessment. By decomposing the evaluation of paper novelty into three structured stages, we constructed an evaluation model, GoAI-CoT-Reviewer, to evaluate the novelty of ideas for assessing, providing feedback, and refining.
3.5.1 Step-by-Step Evaluation Stages
@@
As shown in Figure 3, inspired by the approach of LLaVA-CoT Zhong et al. (2024) and incorporating the thought process of human reviewers, we decompose the evaluation of paper novelty into three structured stages: summary, analysis, and scoring.
@@
GoAI-CoT-Reviewer uses dedicated tags (e.g., <Summary>...</Summary>) to mark each stage, indicating the beginning and end of each phase. These tags help the model maintain clarity throughout the reasoning process. Unlike traditional CoT Wei et al. (2023) reasoning, which allows the model to freely think, our approach promotes structured thinking by first summarizing the input information, then performing a detailed analysis, and finally arriving at a score. To achieve this, we construct a training dataset by progressively generating responses using open review comments from the Openreview platform and then train the evaluation model using supervised fine-tuning. After training, the model autonomously selects these labels as needed, completing all stages within a single process. This structured approach enables the model to independently manage its reasoning process, improving its performance on complex tasks that require multi-step analysis.
3.5.2 Data Preparation and Model Training
@@
We use the Openreview API v2555https://docs.openreview.net/reference/api-v2 to scrape open-source AI conference papers and corresponding review comments submitted on the Openreview platform, including both accepted and rejected papers. To ensure the quality of the papers and review comments to some extent, we scraped papers and reviews from top AI conferences such as ICLR 2022, NeurIPS 2022, and ICLR 2023, with a total of over 35k samples in the training dataset.
@@
Following the approach of the LLAVA-CoT-100K dataset Xu et al. (2025), we use the summary_of_the_paper field from the review comments as the golden answer for the Summary stage, the summary_of_the_paper and strengths_and_weaknesses fields as the golden answers for the Analysis stage, and the technical_novelty_and_significance field as the golden answer for the Scoring stage to construct a supervised fine-tuning dataset. We fine-tune the base model using LoRA to obtain the GoAI-CoT-Reviewer model, which is used for novelty evaluation.
4 Experiments
4.1 Setup
Due to the large number of papers and their lengthy nature, and to reduce costs, we use the GLM-4-Flash API [^6] https://bigmodel.cn/dev/activities/free/glm-4-flash as the backend large language model. The GoAI-CoT-Reviewer evaluation model is built on the Llama 3.1-8B-Instruct model Dubey et al. (2024) and fine-tuned through LoRA. Training is conducted using NVIDIA A100 GPUs, with a LoRA rank set to 8, an input truncation length of 2048, and the number of epochs set to 3.
We use Idea Arena Li et al. (2024), a pairwise evaluation system, to assess the novelty of ideas generated by large language models. Idea Arena utilizes a round-robin competition to calculate the Elo score for each idea generation method. For a given topic, we ask large language model judges to rank the generated ideas based on dimensions such as novelty, significance, clarity, and expected effectiveness. The Elo scores for each method are then calculated using the win-loss-tie records obtained from these evaluations Zheng et al. (2023); Zhao et al. (2024). A detailed description of the evaluation metrics can be found in our supplementary materials.
Method Novelty Significance Effectiveness Clarity Feasibility Average
GoAI (ours) 3.83 3.27 3.6 3.02 3.22 3.388
Chain of Ideas Li et al. (2024) 3.21 3.17 3.77 2.95 3.24 3.268
AI-Scientist Lu et al. (2024) 2.79 2.88 2.58 3.05 2.83 2.826
ResearchAgent Baek et al. (2024) 2.73 2.87 2.67 2.94 2.98 2.838
Vanilla LLM 2.44 2.81 2.38 3.04 2.85 2.704
Table 1: Evaluation results of idea generation.

### v2

# GoAI: Enhancing AI Students’ Learning Paths and Idea Generation via Graph of AI Ideas

@@ Abstract
With the rapid advancement of artificial intelligence technology, AI students are confronted with a significant ”information-to-innovation” gap: they must navigate through the rapidly expanding body of literature, trace the development of a specific research field, and synthesize various techniques into feasible innovative concepts. An additional critical step for students is to identify the necessary prerequisite knowledge and learning paths. Although many approaches based on large language models (LLMs) can summarize the content of papers and trace the development of a field through citations, these methods often overlook the prerequisite knowledge involved in the papers and the rich semantic information embedded in the citation relationships between papers. Such information reveals how methods are interrelated, built upon, extended, or challenged. To address these limitations, we propose GoAI, a tool for constructing educational knowledge graphs from AI research papers that leverages these graphs to plan personalized learning paths and support creative ideation. The nodes in the knowledge graph we have built include papers and the prerequisite knowledge, such as concepts, skills, and tools, that they involve; the edges record the semantic information of citations (e.g., baseline, extension, comparison). When a student queries a specific paper, a beam search-based path search method can trace the current development trends of the field from the queried paper and plan a learning path toward cutting-edge objectives. The integrated Idea Studio guides students to clarify problem statements, compare alternative designs, and provide formative feedback on novelty, clarity, feasibility, and alignment with learning objectives. Test results, including those on learning gains, creativity quality, and cognitive load, indicate that graph-based learning path planning, creativity generation, and structured feedback can enhance students’ ability to summarize frontiers, identify key technologies, and generate feasible concepts.
Introduction
Generating feasible and innovative ideas lies at the heart of AI education. Traditionally, this process requires students to read academic papers in the AI domain, trace research trends, identify and master the prerequisite knowledge embedded within those works, and, based on the prevailing research directions and cutting-edge advancements, formulate novel ideas. However, in the rapidly evolving field of artificial intelligence—characterized by an explosive growth in publications and increasingly interdisciplinary research—students are often overwhelmed by the sheer volume and intricate citation networks of scientific literature. As a result, they must devote significant time and effort to navigate this landscape. The insufficiency of comprehensive literature reviews further exacerbates the difficulty, making it challenging for students to gain a holistic understanding of a topic or to chart a personalized learning pathway from their current knowledge base toward the generation of innovative ideas.
Although LLM-based academic assistants are capable of answering questions, summarizing papers, and even generating novel ideas (Wang et al. 2024b, a; Li et al. 2024), they typically provide only a linear narrative of research development, which presents three fundamental limitations: (i) they overlook pedagogical prerequisites—namely, what foundational knowledge must be acquired first and how the difficulty should progress—rendering the output less practical for direct student use; (ii) they fail to capture the semantic roles of citations, such as whether a cited work serves as a baseline to be surpassed, an extensible foundation, or a critique to be addressed; and (iii) their linear depiction oversimplifies the intricate structure of academic development, which is better represented as a graph where the interconnections among scholarly works form a complex knowledge network. These limitations hinder both learning efficiency and the quality of idea generation. In essence, the intellectual landscape that AI students must navigate is inherently graph-structured: research papers and their corresponding ideas or proposed methods serve as nodes; citation relationships—each carrying functional semantic roles—form the edges; and the concepts and skills embedded in these works create prerequisite chains. A knowledge graph perspective unveils this structure, enabling algorithmic planning of personalized learning pathways and revealing “innovation opportunities”—points where methods may be integrated, extended, or contrasted to inspire creative research directions.
To address these limitations, we introduce GoAI, a novel framework that leverages knowledge graphs and large language models to analyze the relationships among AI research ideas and generate personalized learning pathways for students. First, we construct a research-centric knowledge graph for AI education by extracting paper entities and semantically annotated citation relationships. These semantic citation links encode the functional roles between scholarly works—such as baseline, extension, contrast/ablation, or critique/response—while the paper entities encompass not only the publications themselves but also the methods proposed, concepts involved, and prerequisite knowledge or skills required. Second, given a student’s target topic or research interest, we retrieve the most relevant papers from the graph and employ beam search guided by citation semantics to generate multiple development trajectories, summarizing the evolution of research in the field. The search process prioritizes citation types most conducive to achieving the learning objective (e.g., understanding baselines before exploring extensions) and organizes the resulting papers into a progressive curriculum based on the complexity of their prerequisite concepts. Finally, within the integrated Idea Studio, students synthesize their acquired knowledge into project ideas , receiving formative feedback from a Chain-of-Thought (CoT) reviewer that evaluates the novelty, strengths, and potential limitations of their proposals.
The main contributions of this study are summarized as follows
  • We propose GoAI, a knowledge graph for AI education that is centered around research papers and enriched with metadata such as concepts and skills, while preserving semantically meaningful citation labels. This integrated structure bridges pedagogy and scholarly research, supporting both learning and creative ideation.
  • Building upon GoAI, we introduce a learning path planning method in which large language models dynamically traverse the AI graph to analyze research trends and generate personalized learning trajectories.
  • We design an Idea Studio that assists students in refining their research ideas based on identified trends, and enhances the transparency of the evaluation process through structured assessments.
  • Experimental results demonstrate that the GoAI framework effectively reduces students’ cognitive load, leading to measurable improvements in both learning outcomes and the quality of generated ideas.
Related Works
Knowledge Graph in Education
In the field of education, knowledge graphs have made notable progress in relationship modeling and practical applications, with increasing maturity in large-scale resource integration and evaluation ecosystems (Abu-Salih and Alotaibi 2024; Qu et al. 2024). In terms of graph construction, EDUKG (Zhao et al. 2022), designed for K–12 education, integrates textbooks and educational resources through fine-grained ontologies and sustainable maintenance mechanisms, laying a foundation for unified modeling across interdisciplinary knowledge and resources. Other efforts, such as interactive annotation pipelines leveraging human–AI collaboration (Aytekin and Saygın 2024), have significantly reduced expert annotation workload. Automated extraction and fusion frameworks for multi-source educational knowledge graphs have also been explored and validated in domain-specific contexts (Li et al. 2023). On the application side, early representation learning methods within MOOCs (Pan et al. 2017) and recent enhancements using directed GNNs to capture richer dependencies (Qu, Shang, and Zhang 2024) highlight sustained advancements in learning path planning and curriculum dependency modeling. Knowledge graphs have been integrated with semantic encoders such as GNNs and SBERT to enable interpretable and fine-grained recommendation systems on MOOC platforms (Alatrash et al. 2024). EDUKG has further been employed to provide factual context for LLMs to generate auditable and explainable learning recommendations (Abu-Rasheed, Weber, and Fathi 2024), and graph-based retrieval-augmented generation (GraphRAG) approaches tailored to educational path planning have begun to emerge (Abdelmagied et al. 2025). However, these studies have largely concentrated on the integration of static knowledge, with limited attention to leveraging knowledge graphs for the synthesis of academic literature and the facilitation of student-driven innovation.
AI-based Scientific Discovery
@@ Introduction
Artificial intelligence has become a powerful tool for advancing research processes, reshaping scientific discovery by reducing the time needed for experimental iterations, and assisting researchers in various capacities (Yu, Xu, and Gao 2024; Abramson et al. 2024; de Haan et al. 2024; Tan, Xin, and Wu 2024). The development of large language models, with their advanced natural language processing capabilities and ability to integrate interdisciplinary information, enables more efficient literature analysis and the generation of novel hypotheses (Kumar et al. 2024; Si, Yang, and Hashimoto 2024). Prior research has enhanced novelty by iteratively comparing AI-generated ideas with existing literature (Wang et al. 2024a), while other studies have focused on linking academic knowledge through core papers, expanding understanding via academic graphs and entity-centered knowledge stores (Baek et al. 2024; Wang et al. 2024b; Li et al. 2024). Some studies use multi-agent systems based on LLMs to foster the generation of research ideas in autonomous scientific discovery (Su et al. 2024). In this paper, we adopt a paradigm for generating research ideas based on academic graphs derived from core papers. Unlike previous works, we represent citation relationships between academic papers and summarize development trends through a knowledge graph, rather than through research trend chains. Furthermore, we introduce the concept of constructing a knowledge graph that incorporates semantic information and citation importance, providing richer information for analyzing relationships between papers.
Graph-augmented LLMs
@@
The integration of LLMs with knowledge graphs, which store information in natural language, is an emerging research area aimed at fully utilizing the explicit and updatable external knowledge within structured graphs to enhance the capabilities of large language models (Pan et al. 2024; Huang et al. 2024), particularly in tasks such as reasoning (Luo et al. 2024b, c) and rule completion (Luo et al. 2024a) . To input the structured knowledge from knowledge graphs into large language models, some methods transcribe entity-relation-entity triples from the knowledge graph into natural language text that is understandable by the language model, which is then provided as a prompt to the LLM (Sun et al. 2024), similar to retrieval-augmented generation. Other methods directly input graph information processed by approaches such as Graph Neural Networks (GNNs) into the large language model, utilizing the implicit graph information for enhancement (Huang et al. 2024; Mavromatis and Karypis 2024). In this paper, since the literature used to construct the graph is already in natural language and the relations within the graph contain semantic information in natural language, we adopt the method of transcribing the graph into textual prompts for input into the large language model to achieve the collaboration between the graph and the model.
@@
Figure 1: The framework of GoAI. The framework consists of four stages: (1) Literature Search and Filtering, (2) GoAI Graph Construction, (3) Path Generation, and (4) Path Validation.
Methodology
GoAI Framework
In this section, we present a detailed overview of the GoAI framework. As illustrated in Figure 1, the framework comprises four stages: (1) literature search and filtering, (2) construction of the GoAI knowledge graph, (3) path generation, including generating research trends and learning pathways, and (4) validation of pathway coherence. Given a research topic, GoAI first searches online academic repositories for relevant literature and identifies key reference papers based on their ranking. It then expands the knowledge graph bi-directionally by incorporating papers connected via citation links to these key references. For each paper, the core idea is extracted to reflect the research trajectory of the field, while the embedded AI-specific prerequisite knowledge is gathered to support student learning. Subsequently, an LLM-as-agent analyzes the graph to synthesize the evolving research trends and constructs a personalized, prerequisite-aligned learning path. Based on this trajectory, the model generates “Hint Ideas” to inspire innovation. Finally, a multi-agent evaluation system assesses these Hint Ideas and learning paths, validating the coherence of the proposed pathways through iterative refinement and integration, thereby ensuring the plausibility and academic soundness of the generated research progression and learning paths.
Literature Search and Filtering
@@
We use the Semantic Scholar API [^1] https://www.semanticscholar.org/product/api and Arxiv API [^2] https://info.arxiv.org/help/api/index.html to search for papers in the online paper database. Given an initial research topic, we retrieve the top-ranked papers, which we refer to as key references, forming the foundation for building the graph. Starting from the corresponding key reference papers, GoAI constructs the graph by extending in both forward and backward directions based on citation relationships. The backward extension is derived from the reference list of the key reference, tracing the origins of the paper’s ideas. The forward extension is achieved by exploring which papers in the literature cite the key reference paper, representing the subsequent development of the idea.
@@
Graph Construction
GoAI Graph Overview
The GoAI graph is used to represent the relationships between papers that have citation relationships, including the citing and cited papers, the importance of the citation (measured by the position where the reference appears in the original text), and the semantic information contained in the citation. The general representation of the GoAI graph is a structured <<paper1, citation position, citation semantics, paper2>> quadruple. The citation position and citation semantics together form the relation field in traditional knowledge graphs, which is used for subsequent analysis of the graph. To construct the GoAI graph, it is necessary to separately obtain the papers as entities and the citation positions and citation semantics as relations.
Relation Definition and Extraction
@@
Path Generation
The process of generating paths from the graph is divided into two steps: graph exploration and path generation. Figure 2 illustrates an example workflow of graph exploration and path generation.
Graph Exploration
In the graph, the entities are papers, and the relations are represented as pairs of citation positions and citation semantics. At the beginning of the $D$-th iteration, each path consists of 2D12D-1 quadruples, i.e., $p=\{(e^{d}_{s},r^{d},e^{d}_{o})\}^{2D-1}_{d=1}$, where $e_{s}$ and $e_{o}$ represent the subject and object entities, respectively, and $r$ denotes the specific relation between them. The pairs $(e^{d}_{s},r^{d},e^{d}_{o})$ and $(e^{d+1}_{s},r^{d+1},e^{d+1}_{o})$ are interconnected. The sets of endpoint entities and relations of all paths in the path set $P$ are denoted as $E^{D-1}=\{e^{d-1}_{1},e^{d-1}_{2},\cdots,e^{d-1}_{N}\}$ and $R^{D-1}=\{r^{d-1}_{1},r^{d-1}_{2},\cdots,r^{d-1}_{N}\}$, respectively.
The exploration phase of the $D$-th iteration aims to utilize the LLM to identify the top-N most relevant entities $E^{D}$ from the neighboring entities of the current top-N endpoint entity set $E^{D-1}$, and use $E^{D}$ to extend the top-N paths in the path set $P$. The endpoint entities here can be either the head or tail entities of the path, corresponding to backward and forward extensions, respectively. To address the complexity of processing a large number of neighboring entities using the LLM, inspired by the approach of ToG (Sun et al. 2024), this paper implements a two-step exploration strategy: first, explore the important relations, and then use the selected relations to guide entity exploration.
Relation Exploration The relation exploration is a beam search process with a depth of 1 and a width of $N$, from $E^{D-1}$ to $R^{D}$ . The entire process can be broken down into two steps: search and pruning. We use the GoAI Agent, an LLM-based agent, to automatically complete this process.
@@
  • Search Step Once we obtain the top-N new paths $P$ and the new set of relations $R^{D}$ from the relation exploration, for each relation path $p_{n}\in P$, we can search $(e^{D-1}_{n},r^{D}_{n},*)$ or $(*,r^{D}_{n},e^{D-1}_{n})$ to obtain the candidate entity set $E^{D}_{cand,n}$, where $e^{D-1}_{n}$ and $r_{n}$ represent the endpoint entities and relations of $p_{n}$. We can then aggregate $\{E^{D}_{cand,1},E^{D}_{cand,2},\cdots,E^{D}_{cand,n}\}$ into $E^{D}_{cand}$ and use the endpoint entities $E^{D}_{cand}$ to extend the top-N paths $P$ to $P_{cand}$. For the shown case, $E^{1}_{cand}$ can be represented as {Self-Consistency, CoT, CPO, Diagram of Thought, ControlLLM, $\cdots$}.
@@
Research Tread and Learning Path Generation with paths
Upon obtaining the current exploration path $p$, we prompt the LLM to evaluate the developmental trajectory represented by the papers along this path, yielding a synthesized trend $T$. Concurrently, for each generated path, the model is guided to predict potential future research directions. Following this, we further prompt the LLM to articulate the underlying motivation, novelty, and methodological approach, ultimately producing a Hint Idea $H$. This Hint Idea serves a dual purpose: it can be utilized in the subsequent section by the novelty-evaluation agents to evaluate candidate ideas as a proxy for path plausibility, and it also functions as an important stimulus and reference point for fostering student creativity.
Similarly, once the current path $p$ is determined, the LLM is prompted to extract the AI-related concepts and skills embedded in each paper along the path and to sequence them in ascending order of complexity, thereby constructing a prerequisite-aligned learning path essential for mastering the developmental trajectory of the field.
Path Validation
We construct a multi-agent evaluation system to assess the feasibility of the proposed pathways, encompassing both the research trajectory and the associated learning path. For the learning path, the evaluating agents are prompted to verify the accuracy of the extracted prerequisite knowledge from each paper and to assess the logical sequencing of the learning progression. Revisions to each path are made through a majority voting mechanism among the agents.
For the research trajectory, we assess the plausibility and potential of the direction by evaluating the novelty of the LLM-generated Hint Idea $H$. Given that evaluating the originality of a research idea requires in-depth analysis and nuanced reasoning, we adopt a structured, step-by-step evaluation process. Specifically, we decompose the assessment of idea novelty into three structured stages, and develop a dedicated evaluation model, GoAI-CoT-Reviewer, to perform this task through chain-of-thought reasoning (Wei et al. 2023), enabling a more rigorous and interpretable judgment of innovation.
Step-by-Step Evaluation Stages
@@
As shown in Figure 3, incorporating the thought process of human reviewers, we decompose the evaluation of paper novelty into three structured stages: summary, analysis, and scoring.
@@
GoAI-CoT-Reviewer uses dedicated tags (e.g., <<Summary>><</Summary>>) to mark each stage, indicating the beginning and end of each phase. These tags help the model maintain clarity throughout the reasoning process. Unlike traditional CoT reasoning, which allows the model to freely think, our approach promotes structured thinking by first summarizing the input information, then performing a detailed analysis, and finally arriving at a score. To achieve this, we construct a training dataset by progressively generating responses using open review comments from the OpenReview platform and then train the evaluation model using supervised fine-tuning. After training, the model autonomously selects these labels as needed, completing all stages within a single process. This structured approach enables the model to independently manage its reasoning process, improving its performance on complex tasks that require multi-step analysis.
Data Preparation and Model Training
@@
We use the Openreview API v2444https://docs.openreview.net/reference/api-v2 to scrape open-source AI conference papers and corresponding review comments submitted on the Openreview platform, including both accepted and rejected papers. To ensure the quality of the papers and review comments to some extent, we scraped papers and reviews from top AI conferences such as ICLR 2022-2025, NeurIPS 2022-2025, with a total of over 38k samples in the training dataset.
@@
We use the summary of the paper field from the review comments as the golden answer for the Summary stage, the summary of the paper and strengths and weaknesses fields as the golden answers for the Analysis stage, and the technical novelty and significance field as the golden answer for the Scoring stage to construct a supervised fine-tuning dataset. We fine-tune the base model using LoRA to obtain the GoAI-CoT-Reviewer model, which is used for path validation.
Multi-agent Evaluation
After training the GoAI-CoT-Reviewer, we assemble a multi-agent evaluation system comprising multiple GoAI-CoT-Reviewer agents to assess the novelty of each Hint Idea, thereby providing a reference for determining the plausibility and value of the corresponding research trajectory. Similar to the validation process for learning paths, each GoAI-CoT-Reviewer agent is prompted to generate a qualitative assessment and assign a numerical score ranging from 1 to 10. A threshold score of 5 is employed to distinguish between promising and unpromising ideas, aligning with the acceptance criteria used in peer review. The final judgment on the quality of an idea is then determined through a majority vote among agents.
SUS $\uparrow$ NASA-TLX $\downarrow$ Avg. Idea Score $\uparrow$
Mental Physical Temporal Performance Effort Frustration Avg.
Group 1 (Web) \ 15.33 7.00 16.67 13.67 17.00 13.33 13.83 5.67
Group 2 (LLM) 85.33 8.33 7.67 7.67 8.67 6.67 7.33 7.72 6.33
Group 3 (GoAI) 87.67 5.33 6.33 3.67 8.67 4.67 7.67 6.06 7.33
Table 1: The performance of student groups employing different methods.